OpenStudy (anonymous):

Solve for y: 2x + y = 4, 3x - 2y = -1

OpenStudy (shubhamsrg):

try hit and trial (ain't that difficult,,trust me) ;)

OpenStudy (anonymous):

[[y = -2*x+4]], [[y = (3/2)*x+1/2]] shouldn't you just do it individually?